# Approximating the universal thermal climate index using sparse regression with orthogonal polynomials

### v1

@@ 3 Results and discussion
Figure 7: (a) Comparison of error histograms for the standard UTCI approximation and the sparse regression fit for a degree 10 Legendre polynomial. (b) Cumulative distribution for the absolute errors of the two models.
This suggests that through the use of sparse regression with an orthogonal polynomial basis we have achieved a Fourier-like decomposition of the UTCI Offset in the Legendre basis (instead of the trigonometric one). This has a number of theoretical advantages: due to the orthogonality of the basis functions, the decomposition minimizes the $L_{2}$ distance (least squares) between approximation and function, guaranteeing the best possible polynomial fit for a given model complexity [52]. Additionally, the coefficients are uncorrelated and hierarchically structured, ensuring that lower-order components remain stable as higher-order terms are added—enhancing both interpretability and numerical robustness. Fig. 6(b) shows the spatial distribution of the error between the proposed sparse regression fit (using a degree 10 Legendre polynomial) and the true UTCI Offset values for a fixed relative humidity of 5%. The errors are small and smoothly varying, indicating good agreement across the input space.
Fig. 7(a) presents a comparison of error histograms for both the standard approximation and the sparse model. The sparse model produces a narrower, more sharply peaked distribution centered at zero, highlighting a reduction in error variance and suggesting better generalization. Fig. 7(b) shows the cumulative distribution of absolute approximation errors for two models. The curve for our fit rises more steeply and reaches higher cumulative values at lower error thresholds, indicating that a larger proportion of predictions fall within smaller error margins. This reflects a clear improvement in accuracy: the sparse model not only reduces average error but also minimizes the frequency of large deviations compared to the standard approximation. For example, the frequency of absolute errors larger than $2^{\circ}$C is reduced from 8% to 4%, while the frequency of errors larger than $3^{\circ}$C reduced from 2% to 0.5%. These results confirm that the sparse regression approach can achieve comparable or improved predictive accuracy while maintaining interpretability and model parsimony.
4 Conclusion
This work has addressed key challenges in environmental and climate data modeling, with a focus on the Universal Thermal Climate Index (UTCI) as a complex, physiologically based measure of thermal comfort. We aimed at improving the approximation of the UTCI compared the to literature baseline. Accurately modeling such indices requires interpretable and efficient regression methods. Symbolic regression and grammar-based approaches—particularly when combined with sparse modeling—offer a powerful framework for discovering concise, physically meaningful relationships in data, enabling robust equation discovery in complex environmental systems.
@@ 4 Conclusion
Sparse regression methods, such as LASSO, helped reduce overfitting and improved interpretability. As we have shown, the choice of basis functions is crucial: orthogonal polynomials like Legendre polynomials offer better numerical stability and conditioning than monomials. They enable hierarchical models where higher-order terms don’t affect lower-order estimates, making them especially useful in sparse, interpretable models. Empirical results support these theoretical advantages. Using sparse regression with orthogonal polynomial basis we have:
@@
    Achieved better accuracy (lower loss with same number of parameters) than the state of the art in the literature.
@@
    Found a Pareto front for different model complexities. Loss curves reveal that sparse models with orthogonal bases efficiently populate a Pareto front, balancing complexity and accuracy.
@@
    Determined coefficients consistent over models with different capacities. Coefficient plots for models built on orthogonal bases show the progressive inclusion of higher-order components without disrupting lower-order structure, in contrast to models using simple regression and ordinary polynomials.
@@
    Trained the model over only 20% of the data with successful generalization in testing over the other 80%. The results are robust under bootstrapping.
@@
Additionally, error distributions for better models exhibit sharp, Gaussian-like peaks, reflecting low and consistent residuals. Visualizations of approximation error over a three-dimensional input space confirm this, with better models yielding tightly constrained error magnitudes across the domain. Together, these insights underscore the synergy between sparse modeling, orthogonal representations, and equation discovery methods in building interpretable, efficient models for complex environmental phenomena such as described by UTCI. Future work will extend this framework by integrating grammar-based equation discovery with sparse regression in orthogonal bases to uncover compact, physically consistent representations of different environmental indices and measurements.
The code to reproduce the results is available at:
https://doi.org/10.5281/zenodo.16880383
@@ 4.0.1 Acknowledgements
This publication is supported by the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie Postdoctoral Fellowship Programme, SMASH co-funded under the grant agreement No. 101081355. The operation (SMASH project) is co-funded by the Republic of Slovenia and the European Union from the European Regional Development Fund.
The authors acknowledge the financial support of the Slovenian Research Agency via the Gravity project AI for Science, GC-0001 and of the Slovenian Research And Innovation Agency (research core funding No. P1-0188).

### v2

@@ 3 Results and discussion
Figure 5: (a) Spatial distribution of the UTCI Offset error (approximation minus reference) for the new sparse-model-based approximation at a fixed relative humidity of 5%, showing small, smoothly varying discrepancies. (b) Comparison of error histograms for the standard UTCI approximation and the new approximation based on the tenth-degree Legendre polynomials. (c) Cumulative distributions of the absolute errors of the two approximations.
Table 3 summarizes the most relevant properties of the two approximations. The results show a clear improvement in accuracy: the new approximation not only substantially reduces the average errors (i.e, the mean error, the mean absolute error, and the root mean square error) but also drastically reduces the frequency of large deviations compared to the standard approximation. For example, the frequency of absolute errors larger than $2^{\circ}$C is halved from 8% to 4%, the frequency of errors larger than $3^{\circ}$C reduces from 2% to 0.5%, while the frequency of errors larger than $4^{\circ}$C reduces from 0.3% to 0.01%. These results clearly show the added benefits of the new approximation and confirm that the sparse regression approach can achieve comparable or improved predictive accuracy while maintaining interpretability and model parsimony.
Standard New
approximation approximation
Polynomial degree 6th 10th
Basis functions monomials Legendre
Number of coefficients 210 209
Mean Error $1.7\cdot 10^{-3}$ C (0.35 C) $2.7\cdot 10^{-15}$ C (0.22 C)
Mean Absolute Error 0.81 C (1.33 C) 0.64 C (0.71 C)
Root Mean Square Error 1.17 C (2.77 C) 0.88 C (0.96 C)
Freq. of abs. errors larger than 2C 8.4 % (15.5 %) 4.2 % (5.0 %)
Freq. of abs. errors larger than 3C 2.2 % (6.3 %) 0.50 % (0.60 %)
Freq. of abs. errors larger than 4C 0.34 % (3.8 %) 0.011 % (0.10 %)
Freq. of abs. errors larger than 5C 0.038 % (3.3 %) 0.00096 % (0 %)
Table 3: Comparison of properties of the standard (Bröde et al., 2012) and new polynomial approximations of UTCI Offset function. The values outside of the parentheses reflect the evaluation of the approximations on the full dataset of 104 643104\,643 accurate Offset values provided by Bröde et al. (2012). The values shown in the parentheses reflect the evaluation using the independent dataset of 1000 accurate UTCI values (Bröde, 2021b), which were not used during the development of the new approximation. Both approximations are only valid for the intervals of environmental variables available in the full dataset (Table 1).
We also evaluated the new approximation on the independent dataset of 1000 accurate UTCI values, which were not used during the development of the approximation. This dataset was prepared by the authors of the Bröde et al. (2012) paper, and is freely available on a Zenodo repository (Bröde, 2021b). Similarly to the evaluation of the new approximation on the full dataset, evaluation on the independent dataset shows a substantial reduction of the mean errors and a drastic reduction in the frequency of large errors compared to the standard approximation (Table 3).
Since the new approximation was determined using the full dataset of accurate Offset values (Bröde et al., 2012), it is, same as the standard approximation, only valid for the intervals of environmental variables available in this dataset (Table 1). Using the approximation for conditions outside of these intervals can potentially lead to large errors or unrealistic results and should be avoided.
4 Conclusions
The goal of this study was to develop an improved version of the polynomial approximation – one that would have comparable computational complexity to the existing approximation but would be more robust in terms of numerical stability and substantially more accurate, particularly in reducing the frequency of larger errors. This goal was successfully achieved using sparse regression with an orthogonal polynomial basis.
@@ 4 Conclusions
Sparse regression methods, such as LASSO, helped reduce overfitting and improve interpretability. As we have shown, the choice of basis functions is crucial: orthogonal polynomials like Legendre polynomials offer better numerical stability and conditioning than monomials. They enable hierarchical models where higher-order terms don’t affect lower-order estimates, making them especially useful in sparse, interpretable models. Empirical results support these theoretical advantages.
Using sparse regression with an orthogonal polynomial basis (or sparse orthogonal regression), we have:
@@
    Achieved substantially better accuracy – compared to the standard approximation, the new approximation not only substantially reduces the average errors (i.e, the mean error, the mean absolute error, and the root mean square error) but also drastically reduces the frequency of large errors.
@@
    Retained a comparable computational complexity – the number of coefficients is almost the same for both approximations, meaning the computational complexity is comparable.
@@
    Found a Pareto front for different model complexities – loss curves reveal that sparse models with orthogonal bases efficiently populate a Pareto front, balancing complexity and accuracy.
@@
    Determined coefficients consistent over models with different capacities - coefficient plots for models built on orthogonal bases show the progressive inclusion of higher-order components without disrupting lower-order structure, in contrast to models using simple regression and ordinary polynomials.
@@
    Achieved successful generalization – training the model over only 20% of the data, while testing was performed over the other 80%, highlights successful generalization. The results are also robust under bootstrapping.
  6. (f)
@@
We have also prepared an easy-to-use Python function for the new approximation (please refer to the Code and data availability section on how to obtain the code). The code relies only on basic mathematical operations, which makes it easy to adapt to other programming languages, such as Fortran or C++. We also implemented a check to see if the environmental state falls within the domain of validity of the approximation. If this is not the case, the code produces a warning that the resulting UTCI values could have large errors or be unrealistic.
Funding
This publication is supported by the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie Postdoctoral Fellowship Programme, SMASH co-funded under the grant agreement No. 101081355. The operation (SMASH project) is co-funded by the Republic of Slovenia and the European Union from the European Regional Development Fund. The authors acknowledge the financial support of the Slovenian Research Agency via the Gravity project AI for Science, GC-0001 and of the Slovenian Research And Innovation Agency (research core funding No. P1-0188).
Author Contributions
S.R. - Conceptualization, Data curation, Formal analysis, Investigation, Methodology, Software, Validation, Visualization, Writing (original draft preparation), G.S. - Conceptualization, Resources, Validation, Software, Writing (review and editing), L.T. - Conceptualization, Methodology, Project administration, Supervision, Validation, Writing (review and editing), S.D. - Conceptualization, Funding acquisition, Project administration, Resources, Supervision, Writing (review and editing).
Code and data availability
The Python function code for the calculation of the new approximation as well as all the code needed to reproduce the results of the analysis is published on Zenodo (Roman, 2025a).
Conflict of Interest Statement
The authors declare no conflicts of interest.